VEGAS Team Meeting

GB-137-hub
192.33.117.12##7144
434-817-6443# (telephone)

Attendees: Randy, Adam, Joe, Dave, Sam, Mark, Paul, Marty, Jason, Melinda

Agenda:

  1. Status of BOFs for Modes 2 and 3 (H16k)
  2. Tsys stability
  3. Modes 14/19
  4. fcenter shift
  5. non-constant baselines
  6. Scan to scan integration number variation
  7. astrid scripts to be used for testing vegas (in the context of M&C integration and regression tests)
  8. FTarget values
  9. Subbeam Nod
  10. Status of DIBAS Support
  11. Plans for test time

Meeting notes:

  1. Status of BOFs for Modes 2 and 3 (H16k)
    • BOF rebuild by Hong, and installed for testing/calibration
    • No need to do MMCM tests anymore since last update by Paul
    • High priority tests (w/"spike finder")
  2. Tsys stability
    • Tsys is higher during the first integration (GBTIDL high by a factor of two)
    • Solved by clearing buffers after first integration?
      • No test case required clearing the buffer but looks like the fix should work
        • W/E data should have been affected
      • Sam & Richard to look at fits issues
        • Research if the counts are what is off by factor of two
  3. Modes 14 & 19
    • Version 148 is ready and works in simulation
    • Joe will test
  4. fcenter shift
    • Removed from MR until Science consensus
    • Medium Priority
  5. non-constant baselines:
    • Re-test (Kobelski)
  6. Scan to scan integration number variation
    • Obs noted that w/30 sec integration only got 29 sec. integrations
    • Data needs to be reviewed (Prestage)
    • Actual count may vary within a small window
  7. Astrid scripts to be used for testing vegas
    • simple scripts that exercise basic functions for M&C integration
    • More involved for regression tests - team can help with VEGAS part of regression tests
  8. Solved? IFTarget values - Yup
    • No problem with 3.5 GHz bandwidth
  9. Solved? Subbeam Nod - Yup
    • Recommends shorter integration times (~0.2 sec) yields 5-10% time improvement
    • Notes on VEGAS commissioning wiki page
    • Spectrometer & VEGAS treated the same way, but only looks at VEGAS bank A
  10. Status of DIBAS Support
    • No report other than pointing to github for ADC info
  11. Plans for test time
    • Currently unscheduled request(s):
      • non-constant baselines (Kobelski)
      • Deep integration tests after successful Mode 2/3 tests (X-Q Band) (Frayer/Kobelski)
    • Monday
      • 1pm BOFs for Modes 2 and 3 (H16k) (Brandt/Kobelski)
      • 3pm Version 148 tests (Mode 14 & 19) (Brandt)
    • Tuesday
    • Wednesday
    • Thursday

Other

  1. Please make sure to properly code your timesheets for the current time period and going forward.
    • Software engineers should be using a GBT ops software CM&E account - Account 893252
    • Electronic engineers should be using a GBT ops electronics CM&E account - Account 432510
    • Commissioning activities should be getting charged to hardware support and testing - Account 432510
    • Scientists assisting with shared-risk commissioning should be charging an observer support account
      • 432540 for the most part. Accounts 142020 (GB-based staff) and 142010 (CV-based staff) for when you are helping with overarching observing support (not problem solving, VEGAS specific)

-- MartyBloss - 2014-08-25
Topic revision: r1 - 2014-08-25, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback